# Propensity Score Alignment of Unpaired Multimodal Data

### v1

@@ D.2 Models
In this section we describe experimental details pertaining to the propensity score and VAE (Yang et al., 2021). SCOT (Demetci et al., 2022) and scGLUE (Cao & Gao, 2022) are used according to tutorials and recommended default settings by the authors.
Loss Functions   The propensity score approach minimizes the standard cross-entropy loss for both modalities, as implemented in PyTorch 2.0.1. The VAE includes, in addition to the standard ELBO loss (with parameter $\lambda$ on the KL term), two cross-entropy losses based on classifiers from the latent space: one, weighted by a parameter $\alpha$ to classify $t$ as in the propensity score, and another, weighted by a parameter $\beta$, that classifies which modality the latent point belongs to.
Hyperparameters and Optimization   We use the Adam optimizer with learning rate $0.0001$ and one cycle learning rate scheduler. We follow Yang et al. (2021) and set $\alpha=1$, $\beta=0.1$, but found that $\lambda=10^{-9}$ (compared to $\lambda=10^{-7}$ in Yang et al. (2021)) resulted in better performance. We used batch size 256 in both instances and trained for either 100 epochs (image) or 250 epochs (CITE-seq).
Architecture   For the synthetic image dataset, we use an 5-layer convolutional network (channels $=32,54,128,256,512$) with batch normalization and leaky ReLU activations, with linear heads for classification (propensity score and VAE) and posterior mean and variance estimation (VAE). For the VAE, the decoder consists of convolutional transpose layers that reverse those of the encoder. For the CITE-seq dataset, we use a 5-layer MLP with constant hidden dimension $1024$, with batch normalization and ReLU activations (adapted from the fully connected VAE in Yang et al. (2021)) as both the encoder and VAE decoder. We use the same architecture for both modalities, RNA-seq (as we process the top 200 PCs) and protein.
Optimal Transport   We used POT (Flamary et al., 2021) to solve the entropic OT problem, using the log-sinkhorn solver, with regularization strength $\gamma=0.05$.
@@ D.3 Data
Synthetic Data   We follow the data generating process Section 3 to generate coloured scenes of two simple objects (circles, or squares) in various orientations and with various backgrounds. The position of the objects are encoded in the latent variable $z$, which is perturbed by a do-intervention (setting to a fixed value) randomly sampled for each $t$. Each object has an $x$ and $y$ coordinate, leading to a $4$-dimensional $z$, for which we consider $3$ separate interventions each, leading to $12$ different settings. The modality then corresponds to whether the objects are circular or square, and a fixed transformation of $z$, while the modality-specific noise $U$ controls background distortions. Scenes are generated using a rendering engine from PyGame as $f^{(e)}$. Example images are given in Figure 4.
@@
CITE-seq Data
We also use the CITE-seq dataset from Lance et al. (2022) as a real-world benchmark (obtained from GEO accession GSE194122). These consist of paired RNA-seq and surface level protein measurements, and their cell type annotations over $45$ different cell types. We used scanpy, a standard bioinformatics package, to perform PCA dimension reduction on RNA-seq by taking the first 200 principal components. The protein measurements (134-dimensional) was processed in raw form. For more details, see Lance et al. (2022).

### v2

@@ D.2 Models
In this section we describe experimental details pertaining to the propensity score and VAE [Yang et al., 2021]. SCOT [Demetci et al., 2022] and scGLUE [Cao and Gao, 2022] are used according to tutorials and recommended default settings by the authors.
Loss Functions   The propensity score approach minimizes the standard cross-entropy loss for both modalities. The VAE includes, in addition to the standard ELBO loss (with parameter $\lambda$ on the KL term), two cross-entropy losses based on classifiers from the latent space: one, weighted by a parameter $\alpha$ to classify $t$ as in the propensity score, and another, weighted by a parameter $\beta$, that classifies which modality the latent point belongs to.
Hyperparameters and Optimization   We use the Adam optimizer with learning rate $0.0001$ and one cycle learning rate scheduler. We follow Yang et al. [2021] and set $\alpha=1$, $\beta=0.1$, but found that $\lambda=10^{-9}$ (compared to $\lambda=10^{-7}$ in Yang et al. [2021]) resulted in better performance. We used batch size 256 in both instances and trained for either 100 epochs (image) or 250 epochs (CITE-seq).
For the VAE and classifiers of experiment 3, we use an Adam optimizer with learning 0.001 and weight decay 0.001 and max epoch of 100 (PerturbSeq) and 250 (single cell images) using batch sizes of 256 and 2048 correspondingly. We follow similar settings as Yang et al. [2021] and implement $\alpha=1$ with $\lambda=10^{-9}$, and since we do not have matched data, $\beta=0$. For the cross-modal prediction models in experiment 3, we use Stochastic Gradient Descent optimizer with learning rate 0.001 and weight decay 0.001 with max epochs 250 and batch size 256. We implement early stopping with delay of 50 epochs which we then checkpoint the last model to use for downstream tasks
Architecture   For the synthetic image dataset, we use an 5-layer convolutional network (channels $=32,54,128,256,512$) with batch normalization and leaky ReLU activations, with linear heads for classification (propensity score and VAE) and posterior mean and variance estimation (VAE). For the VAE, the decoder consists of convolutional transpose layers that reverse those of the encoder.
For the CITE-seq dataset, we use a 5-layer MLP with constant hidden dimension $1024$, with batch normalization and ReLU activations (adapted from the fully connected VAE in Yang et al. [2021]) as both the encoder and VAE decoder. We use the same architecture for both modalities, RNA-seq (as we process the top 200 PCs) and protein.
For the PerturbSeq classifier encoder, we use a 2-layer MLP architecture. Each layer consists of a linear layer with an output feature dimension of 64, followed by Rectified Linear Unit (ReLU) activation, Batch Normalization, and dropout (p=0.1). A final layer with Leaky ReLU activation that brings dimensionality to 128 before feeding into a linear classification head with an output feature dimension of 25.
For the single-cell image encoder classifier, we use a proprietary Masked Autoencoder [Kraus et al., 2023] to generate 1024-dimensional embeddings. Subsequently, a 2-layer MLP is trained on these embeddings. Each MLP layer has a linear layer, Batch Normalization, and Leaky ReLU activation. The output feature dimensions of the linear layers are 512 and 256, respectively, and the latent dimension remains at 1024 before entering a linear classification head with an output feature dimension of 25.
Optimal Transport   We used POT [Flamary et al., 2021] to solve the entropic OT problem, using the log-sinkhorn solver, with regularization strength $\gamma=0.05$.
@@ D.3 Data
Synthetic Data   We follow the data generating process Eq. 1 to generate coloured scenes of two simple objects (circles, or squares) in various orientations and with various backgrounds. The position of the objects are encoded in the latent variable $z$, which is perturbed by a do-intervention (setting to a fixed value) randomly sampled for each $t$. Each object has an $x$ and $y$ coordinate, leading to a $4$-dimensional $z$, for which we consider $3$ separate interventions each, leading to $12$ different settings. The modality then corresponds to whether the objects are circular or square, and a fixed transformation of $z$, while the modality-specific noise $U$ controls background distortions. Scenes are generated using a rendering engine from PyGame as $f^{(e)}$. Example images are given in Fig. 4.
@@
CITE-seq Data   We also use the CITE-seq dataset from Lance et al. [2022] as a real-world benchmark (obtained from GEO accession GSE194122). These consist of paired RNA-seq and surface level protein measurements, and their cell type annotations over $45$ different cell types. We used scanpy, a standard bioinformatics package, to perform PCA dimension reduction on RNA-seq by taking the first 200 principal components. The protein measurements (134-dimensional) was processed in raw form. For more details, see Lance et al. [2022].
PerturbSeq and Single Cell Image Data   We collect single-cell PerturbSeq data (200 genes) and single-cell painting images in HUVEC cells with 24 gene perturbations and a control perturbation, resulting in 25 labels for matching across both modalities. The target gene perturbations are selected based on the 24 genes with the highest number of cells affected by the CRISPR guide RNAs targeting those genes. The PerturbSeq data is filtered to include the top 200 genes with the highest mean count, then normalized and log-transformed. The single-cell painting images are derived from multi-cell images, with each single-cell nucleus centered within a 32x32 pixel box. We use scVI Lopez et al. [2018] to embed the raw PerturbSeq counts into a 128-dimensional space before training the gene expression classifier. Similarly, we train our image classifier using 1024-dimensional embeddings obtained from a pre-trained Masked Autoencoder Kraus et al. [2023], He et al. [2022]. Following matching, we perform cross-modality translation from the single-cell embeddings to the transformed gene expression counts.
D.4 Supplementary Results
In Distribution Out of Distribution
Method Wasserstein-1 ($\downarrow$) L2 Norm ($\downarrow$) Wasserstein-1 ($\downarrow$) L2 Norm ($\downarrow$)
Med (Q1, Q3) Med (Q1, Q3)
PS+OT 4.199 3.280 5.394 7.219
(4.173, 4.226) (3.267, 3.284)
VAE+OT 4.339 3.490 5.629 7.444
(4.314, 4.348) (3.486, 3.495)
Random 4.499 3.826 6.239 7.793
(4.478, 4.525) (3.823, 3.828)
Table 3: Wasserstein-1 and L2 norm distance values for PerturbSeq and single cell images experiments where distance is evaluated between cross-modal predictions and actual gene expression values.